# Forecasting GPU Performance for Deep Learning Training and Inference

### v2

@@ Abstract.
Deep learning kernels exhibit a high level of predictable memory accesses and compute patterns, making GPU’s architecture well-suited for their execution. Moreover, software and runtime system for GPUs further enable optimizations that aim to better utilize the stream multiprocessors, on-chip bandwidth, multiple levels of cache hierarchy, and off-chip high-bandwidth memory. In the context of deep learning, the entire space of models and GPUs is constantly evolving, as newer models emerge with simultaneous upgrades to the device. However, access to newer GPUs is often limited, raising important questions about the performance of new model architectures on existing GPUs, existing models on new GPUs, and new model architectures on new GPUs. To address these questions, we introduce NeuSight, a forecasting framework to predict the performance of a diverse range of deep learning models, for both training and inference, on unseen GPUs, without requiring actual execution of the target model on the target GPU. The framework leverages both GPU hardware behavior and software library optimizations to estimate the end-to-end performance of these models. We observe that prior works in this area (Isaev et al., 2023; Li et al., 2023; Yu et al., 2021) suffer from high absolute error percentages when forecasting performance on unseen models and new GPUs, as they attempt to model the complex task of predicting the latency of a deep learning kernel on a GPU directly using a machine learning approach. Instead, with NeuSight, we decompose the prediction into smaller problems, while bounding the prediction through fundamental performance laws. NeuSight decomposes a single deep learning kernel prediction into smaller working sets called tiles, which are executed independently on the GPU. Tile-granularity predictions are determined using a machine learning approach and aggregated to estimate the end-to-end latency. As such, NeuSight outperforms prior work across a variety of deep learning workloads and the most up-to-date GPUs. It reduces the percentage error from 121.4% and 30.8% to 2.3% in predicting the latency of GPT3 model for training and inference on H100, in comparison to state of the art prior works (Yu et al., 2021; Li et al., 2023), respectively, where GPT3 and H100 were not used to train any framework.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlands.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsisbn: 979-8-4007-0698-1/25/03doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Machine learningccs: Computer systems organization Parallel architectures
@@ 1. Introduction
Prior works in the domain of modeling GPU performance take an analytical model, a linear regression-based model, or a neural network approach (Yu et al., 2021; Li et al., 2023; Isaev et al., 2023). We empirically observe that these solutions exhibit high percentage errors for newer models and GPUs, for both inference and training. They are limited for the proposed use case because they either: (1) use a rough analytical model without information about detailed GPU execution (hardware or software) (2) require execution on a specific GPU to extrapolate latency for newer GPUs (Li et al., 2023; Isaev et al., 2023), or (3) are unable to capture the necessary architectural information of the GPUs (Yu et al., 2021). Additionally, these prior works exhibit high percentage error for newer models and unseen GPUs because they aim to directly model the overall latency of a kernel through machine learning. However, GPU execution, even of a single kernel, is a complex conflation of hardware execution across SMs and software optimizations, which often cannot be captured by a machine learning model trying to determine the latency of the kernel only using certain high-level features such as effective memory bandwidth, compute utilization, and peak flop.
Thus, we develop NeuSight which digresses from prior works as it leverages the insight that popular GPU libraries decompose deep learning kernels into multiple smaller working sets. These small working sets, or tiles, are then dispatched and executed on the GPU independently across the SMs. Based on this observation we partition the end-to-end latency prediction into regular and more manageable sub-problems, predicting the latency at tile-granularity. However, note that the GPU execution is governed by a few performance laws - such as the performance cannot exceed the peak flops or peak memory bandwidth. We capture the above two aspects in our prediction and only use a machine learning model to determine the utilization of the device based on the DNN kernel and GPU properties. The output of the MLP, i.e., utilization of the device, is subsequently used to predict the tile, operator, and the model level latency. Overall, the goal of this entire framework is to distill down to the properties of GPU execution, use machine learning to capture the runtime non-linear behavior, bound the results by fundamental performance laws, and gather the predictions for end-to-end latency predictions. This approach ensures robustness in unseen scenarios, such as new GPUs and models. To ensure prediction for large models that require multiple GPUs, we extend our approach to predict the latency of distributed execution within a server. To do so, we predict the latency of communication operators and integrate these with the per-device execution latencies determined by NeuSight to calculate the overall distributed performance.
@@ 2.2. Deep Learning Execution
Per-device execution. Machine learning practitioners use frameworks such as PyTorch and TensorFlow (PyTorch, 2024b; Google, 2024) to represent and deploy a model as a set of layers. The dataflow of the kernels is managed by the framework and dispatched to the GPU. The corresponding device library then executes each kernel and returns the results to the framework. For example, NVIDIA GPUs utilize the CuDNN library for kernel execution, abstracting the underlying hardware complexity from the user. In this paper, similar to prior works, we assume kernels execute sequentially on the device, which is also empirically validated for the mainstream machine learning frameworks (Zhu et al., 2020). Thus, the predicted latency of each kernel is aggregated based on the dataflow of the model that executes on each device to determine the per-GPU execution latency.
@@
Figure 2. Prediction error of prior works on BMM operator, reported in percentage error. Out of distribution dimensions and GPUs are highlighted. For these results, we trained Habitat and Li et al. models only up to V100, excluding any A100s, H100, and L4.
@@ 3. Motivation
There is a growing need to predict GPU performance for both existing and upcoming models for a variety of reasons: (a) comparing performance across various devices, (b) utilizing estimates to identify GPUs that meet the performance requirements, and (c) evaluating the feasibility of deploying new model architectures on either existing or new GPUs. We empirically validate prior works that aim to predict the performance of GPUs by employing data-driven methods and observe that do not generalize well to new GPUs or new models. These drive the insights for our framework, NeuSight.
In this section, for motivation, we focus on a single kernel type Batched Matrix Multiplication (BMM), commonly found in DNNs. We present end-to-end comparisons (training and inference) in Section 6. We observe that even predicting performance of a single kernel on a GPU requires a careful consideration of hardware architecture and software optimization. This insight is based on empirical validation of prior works that employ machine learning to determine the operator latency.
3.1. Predicting Performance of Batched Matrix Multiplication Using Prior Works
@@ 3.3. Other Related Works
Machine Learning based approaches. Due to the large number of compute lanes in GPUs, they often observe linear performance trends for compute intensive kernels, and hence many prior works leverage regression models to predict the performance of a GPU architecture (Jia et al., 2012; Wu et al., 2015; Li et al., 2023). Some works require extracting low-level hardware features, such as hardware counter values, to make predictions. However, these values are unavailable unless the device is accessible (O’Neal et al., 2017; Punniyamurthy et al., 2017; Wang et al., 2019). Such approaches cannot capture the complex intricacies of hardware features and software optimizations such as tiling for deep learning and do not support cross-generational performance prediction with GPUs due to the lack of hardware. Other works (Yang et al., 2022) propose graph neural networks to predict resource consumption (GPU utilization, memory utilization, and network TX/RX throughput) in distributed deep learning. However, this approach necessitates target GPUs for transfer training.
@@ 4. NeuSight Forecasting
NeuSight forecasts the latency of executing a model on multiple GPUs in a server in three steps. The first and main step forecasts the latency of a single kernel that executes on the GPU atomically. The kernel leverages software and hardware optimizations for GPUs, namely tiling, parallel execution across SMs, on-chip caching, and HBM bandwidth. NeuSight uses architectural details such as memory size, memory bandwidth, number of SMs, L2 cache size, and peak FLOPS to predict the kernel’s latency through a machine learning approach. Our kernel estimation is inspired by how deep learning kernels are disassembled into multiple smaller tiles and mapped onto GPU hardware for execution. This approach contrasts with prior works that rely solely on compute or memory utilization to predict latency. Section 4.1 showcases the execution of a few representative DNN kernels on the GPU. NeuSight supports tensor operators such as General Matrix Multiplication (GEMM), element-wise operators such as addition and ReLU, and reduction-based operators such as Softmax and layer normalization. In the second step, the per-operator latency is aggregated to predict the performance of the DNN workload based on its dataflow graph. In the third step, detailed in Section 5.1, the performance of per-GPU execution is combined with predicted network latencies to determine the performance of a DNN executing across multiple GPUs in a server.
@@ 4.2. Kernel-wise Prediction
Tile-granularity prediction. For each DNN kernel we break it down into multiple identical tiles, instead of directly predicting the latency of the kernel. This insight is motivated from prior works (Yu et al., 2021), where a machine learning model is unable to capture the necessary hardware level and software optimizations on out of distribution dimensions and GPUs. This implies that there are more intricate execution patterns although consistent across the GPUs, but are not easily captured by machine learning models. Thus, we use the machine learning model to make a prediction per tile, and assemble these predictions to determine the latency of the whole operator. This decomposition simplifies the problem into more manageable sub-problems, allowing the machine learning model to be trained on smaller and more straightforward tasks. For example, the typical tile dimensions used by GEMM library ranges from 32 to 256, which is still much smaller than the dimensions of GEMMs in modern DNN workloads. To compute the latency of the entire kernel from the predictions for each tile, we employ the following:
@@ 5. NeuSight Workflow
NeuSight also takes as input the distribution strategy, data parallel width, pipeline parallel depth and schedule, or tensor model parallel width. It is important to highlight that the focus of NeuSight is to effectively predict the performance of DNNs on new or unseen GPUs. Predicting the performance of each kernel is the complex and unsolved aspect tackled by NeuSight, as it involves addressing the non-linearity of GPU software optimizations and hardware execution. For completeness, we provide a solution that can forecast the latency of distributed execution on a multi-GPU server. The goal of this forecasting is to enable researchers to determine latency of a relatively larger model that often cannot fit on a single device due to memory constraints. This allows NeuSight to provide accurate and reliable predictions for the performance of the compute/memory intensity of the DNN while avoiding the variability introduced by complex multi-server network operations. There are prior works (Wang et al., 2023; Rashidi et al., 2020) that already model complex network topologies, which researchers can leverage with NeuSight.
@@ 5.1. Forecasting for Distributed Execution
Pipeline parallel. Pipeline parallel requires forward and backward pass micro-batches to be interleaved to achieve high utilization as per the schedules proposed by prior works (Huang et al., 2019; Narayanan et al., 2019, 2021). For NeuSight, the user provides the pipeline parallel schedule and the framework stitches the execution across various devices to determine the overall latency of the model executing on a server. Currently, NeuSight inserts the necessary pipeline bubbles between the forward and backward pass based on GPipe (Huang et al., 2019) schedule, however can be easily extended to other schedules. NeuSight estimates the latency of these pipeline bubbles based on the size of the microbatch, the number of GPUs, and the latency of send-receive operations.
@@ 7. Conclusion
In this work, we present a novel framework, NeuSight, that forecasts the performance of deep learning execution on GPU servers. In developing NeuSight, we leverage the execution deep learning kernel on a GPU architecture such as tiled execution and use this insight to decompose the prediction problem to smaller sub-problems, i.e., per-tile prediction. For each tile, we use machine learning models to capture the complex non-linear relationship between achieved throughput and input features, and impose performance laws to bound the predicted latency. Results on a variety of models and GPUs, show that NeuSight can provide better predictions than prior works in the area. We open-sourced NeuSight to benefit broader research community.  [^1] NeuSight is available at github.com/sitar-lab/NeuSight

### v3

@@ Abstract.
Deep learning kernels exhibit a high level of predictable memory accesses and compute patterns, making GPU’s architecture well-suited for their execution. Moreover, software and runtime system for GPUs further enable optimizations that aim to better utilize the stream multiprocessors, on-chip bandwidth, multiple levels of cache hierarchy, and off-chip high-bandwidth memory. In the context of deep learning, the entire space of models and GPUs is constantly evolving, as newer models emerge with simultaneous upgrades to the device. However, access to newer GPUs is often limited, raising important questions about the performance of new model architectures on existing GPUs, existing models on new GPUs, and new model architectures on new GPUs. To address these questions, we introduce NeuSight, a forecasting framework to predict the performance of a diverse range of deep learning models, for both training and inference, on unseen GPUs, without requiring actual execution of the target model on the target GPU. The framework leverages both GPU hardware behavior and software library optimizations to estimate the end-to-end performance of these models. We observe that prior work in this area suffers from high absolute error percentages when forecasting performance on unseen models and new GPUs, as they attempt to model the complex task of predicting the latency of a deep learning kernel on a GPU directly using a machine learning approach. Instead, with NeuSight, we decompose the prediction into smaller problems, while bounding the prediction through fundamental performance laws. NeuSight decomposes a single deep learning kernel prediction into smaller working sets called tiles, which are executed independently on the GPU. Tile-granularity predictions are determined using a machine learning approach and aggregated to estimate the end-to-end latency. As such, NeuSight outperforms prior work across a variety of deep learning workloads and the most up-to-date GPUs. It reduces the percentage error from 121.4% and 30.8% to 2.3% in predicting the latency of GPT3 model for training and inference on H100, in comparison to state-of-the-art prior work, respectively, where GPT3 and H100 were not used to train any framework.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1; March 30–April 3, 2025; Rotterdam, Netherlands.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 1 (ASPLOS ’25), March 30–April 3, 2025, Rotterdam, Netherlandsisbn: 979-8-4007-0698-1/25/03doi: 10.1145/3669940.3707265ccs: Computing methodologies Machine learningccs: Computer systems organization Parallel architectures
@@ 1. Introduction
Prior work in the domain of modeling GPU performance takes an analytical model, a linear regression-based model, or a neural network approach (Yu et al., 2021; Li et al., 2023; Isaev et al., 2023). We empirically observe that these solutions exhibit high percentage errors for newer models and GPUs, for both inference and training. They are limited for the proposed use case because they either: (1) use a rough analytical model without information about detailed GPU execution (hardware or software) (2) require execution on a specific GPU to extrapolate latency for newer GPUs (Li et al., 2023; Isaev et al., 2023), or (3) are unable to capture the necessary architectural information of the GPUs (Yu et al., 2021). Additionally, prior work exhibits high percentage errors for newer models and unseen GPUs because they directly model the overall kernel latency using machine learning. However, GPU execution, even of a single kernel, is a complex conflation of hardware execution across SMs and software optimizations, which often cannot be captured by a machine learning model trying to determine the latency of the kernel only using certain high-level features such as effective memory bandwidth, compute utilization, and peak flop.
Thus, we develop NeuSight which digresses from prior work as it leverages the insight that popular GPU libraries decompose deep learning kernels into multiple smaller working sets. These small working sets, or tiles, are then dispatched and executed on the GPU independently across the SMs. Based on this observation we partition the end-to-end latency prediction into regular and more manageable sub-problems, predicting the latency at tile-granularity. However, note that the GPU execution is governed by a few performance laws - such as the performance cannot exceed the peak flops or peak memory bandwidth. We capture the above two aspects in our prediction and only use a machine learning model to determine the utilization of the device based on the DNN kernel and GPU properties. The output of the MLP, i.e., utilization of the device, is subsequently used to predict the tile, operator, and the model level latency. Overall, the goal of this entire framework is to distill down to the properties of GPU execution, use machine learning to capture the runtime non-linear behavior, bound the results by fundamental performance laws, and gather the predictions for end-to-end latency predictions. This approach ensures robustness in unseen scenarios, such as new GPUs and models. To ensure prediction for large models that require multiple GPUs, we extend our approach to predict the latency of distributed execution within a server. To do so, we predict the latency of communication operators and integrate these with the per-device execution latencies determined by NeuSight to calculate the overall distributed performance.
@@ 2.2. Deep Learning Execution
Per-device execution. Machine learning practitioners use frameworks such as PyTorch and TensorFlow (PyTorch, 2024b; Google, 2024) to represent and deploy a model as a set of layers. The dataflow of the kernels is managed by the framework and dispatched to the GPU. The corresponding device library then executes each kernel and returns the results to the framework. For example, NVIDIA GPUs utilize the CuDNN library for kernel execution, abstracting the underlying hardware complexity from the user. In this paper, similar to prior work, we assume kernels execute sequentially on the device, which is also empirically validated for the mainstream machine learning frameworks (Zhu et al., 2020). Thus, the predicted latency of each kernel is aggregated based on the dataflow of the model that executes on each device to determine the per-GPU execution latency.
@@
Figure 2. Prediction error of prior work on BMM operator, reported in percentage error. Out of distribution dimensions and GPUs are highlighted. For these results, we trained Habitat and Li et al. models only up to V100, excluding any A100s, H100, and L4.
@@ 3. Motivation
There is a growing need to predict GPU performance for both existing and upcoming models for a variety of reasons: (a) comparing performance across various devices, (b) utilizing estimates to identify GPUs that meet the performance requirements, and (c) evaluating the feasibility of deploying new model architectures on either existing or new GPUs. We empirically validate prior work aimed at predicting the performance of GPUs by employing data-driven methods and observe that do not generalize well to new GPUs or new models. These drive the insights for our framework, NeuSight.
In this section, for motivation, we focus on a single kernel type Batched Matrix Multiplication (BMM), commonly found in DNNs. We present end-to-end comparisons (training and inference) in Section 6. We observe that even predicting performance of a single kernel on a GPU requires a careful consideration of hardware architecture and software optimization. This insight is based on empirical validation of prior work employing machine learning to predict operator latency.
3.1. Predicting Performance of Batched Matrix Multiplication Using prior work
@@ 3.3. Other Related Works
Machine Learning based approaches. Due to the large number of compute lanes in GPUs, they often observe linear performance trends for compute intensive kernels, and hence many prior work leverage regression models to predict the performance of a GPU architecture (Jia et al., 2012; Wu et al., 2015; Li et al., 2023). Some works require extracting low-level hardware features, such as hardware counter values, to make predictions. However, these values are unavailable unless the device is accessible (O’Neal et al., 2017; Punniyamurthy et al., 2017; Wang et al., 2019). Such approaches cannot capture the complex intricacies of hardware features and software optimizations such as tiling for deep learning and do not support cross-generational performance prediction with GPUs due to the lack of hardware. Other works (Yang et al., 2022) propose graph neural networks to predict resource consumption (GPU utilization, memory utilization, and network TX/RX throughput) in distributed deep learning. However, this approach necessitates target GPUs for transfer training.
@@ 4. NeuSight Forecasting
NeuSight forecasts the latency of executing a model on multiple GPUs in a server in three steps. The first and main step forecasts the latency of a single kernel that executes on the GPU atomically. The kernel leverages software and hardware optimizations for GPUs, namely tiling, parallel execution across SMs, on-chip caching, and HBM bandwidth. NeuSight uses architectural details such as memory size, memory bandwidth, number of SMs, L2 cache size, and peak FLOPS to predict the kernel’s latency through a machine learning approach. Our kernel estimation is inspired by how deep learning kernels are disassembled into multiple smaller tiles and mapped onto GPU hardware for execution. This approach differs from prior work, which relies solely on compute or memory utilization to predict latency. Section 4.1 showcases the execution of a few representative DNN kernels on the GPU. NeuSight supports tensor operators such as General Matrix Multiplication (GEMM), element-wise operators such as addition and ReLU, and reduction-based operators such as Softmax and layer normalization. In the second step, the per-operator latency is aggregated to predict the performance of the DNN workload based on its dataflow graph. In the third step, detailed in Section 5.1, the performance of per-GPU execution is combined with predicted network latencies to determine the performance of a DNN executing across multiple GPUs in a server.
@@ 4.2. Kernel-wise Prediction
Tile-granularity prediction. For each DNN kernel we break it down into multiple identical tiles, instead of directly predicting the latency of the kernel. This insight is motivated from prior work (Yu et al., 2021), where a machine learning model is unable to capture the necessary hardware level and software optimizations on out of distribution dimensions and GPUs. This implies that there are more intricate execution patterns although consistent across the GPUs, but are not easily captured by machine learning models. Thus, we use the machine learning model to make a prediction per tile, and assemble these predictions to determine the latency of the whole operator. This decomposition simplifies the problem into more manageable sub-problems, allowing the machine learning model to be trained on smaller and more straightforward tasks. For example, the typical tile dimensions used by GEMM library ranges from 32 to 256, which is still much smaller than the dimensions of GEMMs in modern DNN workloads. To compute the latency of the entire kernel from the predictions for each tile, we employ the following:
@@ 5. NeuSight Workflow
NeuSight also takes as input the distribution strategy, data parallel width, pipeline parallel depth and schedule, or tensor model parallel width. It is important to highlight that the focus of NeuSight is to effectively predict the performance of DNNs on new or unseen GPUs. Predicting the performance of each kernel is the complex and unsolved aspect tackled by NeuSight, as it involves addressing the non-linearity of GPU software optimizations and hardware execution. For completeness, we provide a solution that can forecast the latency of distributed execution on a multi-GPU server. The goal of this forecasting is to enable researchers to determine latency of a relatively larger model that often cannot fit on a single device due to memory constraints. This allows NeuSight to provide accurate and reliable predictions for the performance of the compute/memory intensity of the DNN while avoiding the variability introduced by complex multi-server network operations. ”Prior work (Wang et al., 2023; Rashidi et al., 2020) already models complex network topologies, which researchers can leverage alongside NeuSight.”
@@ 5.1. Forecasting for Distributed Execution
Pipeline parallel. Pipeline parallel requires forward and backward pass micro-batches to be interleaved to achieve high utilization as per the schedules proposed by prior work (Huang et al., 2019; Narayanan et al., 2019, 2021). For NeuSight, the user provides the pipeline parallel schedule and the framework stitches the execution across various devices to determine the overall latency of the model executing on a server. Currently, NeuSight inserts the necessary pipeline bubbles between the forward and backward pass based on GPipe (Huang et al., 2019) schedule, however can be easily extended to other schedules. NeuSight estimates the latency of these pipeline bubbles based on the size of the microbatch, the number of GPUs, and the latency of send-receive operations.
@@ 7. Conclusion
In this work, we present a novel framework, NeuSight, that forecasts the performance of deep learning execution on GPU servers. In developing NeuSight, we leverage the execution deep learning kernel on a GPU architecture such as tiled execution and use this insight to decompose the prediction problem to smaller sub-problems, i.e., per-tile prediction. For each tile, we use machine learning models to capture the complex non-linear relationship between achieved throughput and input features, and impose performance laws to bound the predicted latency. Results on a variety of models and GPUs, show that NeuSight can provide better predictions than prior work in the area. We open-sourced NeuSight to benefit broader research community.  [^1] NeuSight is available at github.com/sitar-lab/NeuSight